Holding period gain (Show formula)

Suppose you purchased 40 shares of XYZ stock for SAR 350.00 on February 1.

You sell the 40 shares of stock on October 1 of the same year for 672.40.

No dividends were paid during the year.

1. Calculate the holding period gain

2. Calculate the holding period return
Answer #1

Solution :

1. Calculation of Holding period gain :

The formula for calculating the Holding period gain is

= Total Sale price of stock – Total Purchase price of stock

As per the information given in the question is                                      

No. of shares of XYZ stock purchased = 40 ; Purchase price per share = SAR 350.00

Thus the

Total purchase price of stock = No. of shares of XYZ stock purchased * Purchase price per share

= 40 * SAR 350.00

= SAR 14,000

Thus Total purchase price of stock = SAR 14,000

No. of shares of XYZ stock sold = 40 ; Sale price per share = SAR 672.40

Thus the

Total Sale price of stock = No. of shares of XYZ stock sold * Sale price per share

= 40 * SAR 672.40

= SAR 26,896

Thus Total Sale price of stock = SAR 26,896

Thus the Holding period gain = Total Sale price of stock - Total purchase price of stock

= SAR 26,896 – SAR 14,000

= SAR 12,896

Holding period gain = SAR 12,896

2.Calculation of Holding Period Return:

Holding period return is the return earned on an investment for the period during which it is held.

The formula for calculating the holding period return is

= ( Total sale price + Dividends earned during the holding period – Total Purchase Price ) / Total Purchase Price

As per the information given in the question is                                      

Total sale price = SAR 26,896 ; Dividends earned during the holding period = 0 ;

Total Purchase Price = SAR 14,000

Applying the above information in the formula we have

= ( 26,896 – 14,000 ) / 14,000

= 12,896 / 14,000

= 0.921143

= 92.1143 %

= 92.11 % ( When rounded off to two decimal places )

Thus the holding period return = 92.11 %